A shopkeeper purchased 200 bulbs for Rs. 10 each. However, 5 bulbs were fused and had to be thrown away. The remaining were sold at Rs.12 each. What will be the percentage profit?
Answer by mananth(16946)

200 bulbs at rs 10 = 2000
5 bulbs defective
195 bulbs sold at 12
SP= 195*12=2340
profit= 2340-2000=340
profit percent = profit/CP * 100
=340/2000 * 100
=17%